# Towards Accurate and Interpretable Neuroblastoma Diagnosis via Contrastive Multi-scale Pathological Image Analysis

### v1

@@ 3.5 WSI Voting Classification Mechanism
Figure 2: The architecture of the proposed WSI voting mechanism. The original pathological image is standardized, cut, and stacked to generate a sequence of slices. After feature classification of the slices using an SVM classifier, weights are dynamically assigned based on classification confidence, ultimately implementing a weighted soft voting strategy for decision.
To accurately classify the entire WSI and assist in pathological diagnosis, we design a heuristic soft voting classification process based on clinical knowledge rules, as illustrated in Fig. 2. This process is grounded in the classification results of individual patches and incorporates global feature information to ultimately determine the overall classification label of the WSI. Unlike traditional hard voting methods, the soft voting mechanism we employ flexibly assigns weights based on the varying importance of slices, yielding more robust classification. Furthermore, to more finely discern specific tissue structures within slices, such as neuropil or Schwannian stroma, we train an SVM model to identify tissue components accurately. Below, we will elaborate on the voting mechanism’s design and the SVM model’s training and application.
@@ 3.5.1 Voting Mechanism Design
The implementation of WSI classification through a hard voting mechanism is straightforward; it treats the classification results of all patches as equally weighted votes, determining the final classification based on the majority rule. However, the precision of the hard voting method is relatively low. To achieve a more precise classification, we incorporated prior knowledge and designed a soft voting approach for WSI classification. According to principles of pathological analysis, if a patch contains regions with neuropil, it suggests the presence of cell nests, indicating that this patch is likely to belong to a malignant category, specifically ganglioneuroblastoma (GNB) or NB. Conversely, if a patch contains regions of Schwannian stroma, it suggests the absence of cell nests, leaning more towards ganglioneuroma (GN). Furthermore, if Schwannian stroma constitutes 50% or more of the patch, the WSI tends towards GNB or GN; otherwise, it leans towards NB.
Based on this knowledge, we can derive the voting weights for all patches within the WSI, treating each patch’s weight as a displacement vector $\vec{w}$ in the classification space, where the direction corresponds to the category set and the length represents the weight size. By integrating the voting weights of all patches with their classification results, we arrive at the final WSI classification ${\mathrm{Label}}_{\mathrm{WSI}}$:
$P\in RN\times C,\qquad P\in R^{N\times C},$
$W\in RN\times C,\qquad W\in R^{N\times C},$
@@
$\hat{\mathbf{p}}=\frac{\mathbf{s}}{\alpha}=\frac{1}{\alpha}\begin{bmatrix}\sum\limits_{i=1}^{N}w_{i,1}p_{i,1}\\ \sum\limits_{i=1}^{N}w_{i,2}p_{i,2}\\ \vdots\\ \sum\limits_{i=1}^{N}w_{i,C}p_{i,C}\end{bmatrix},$ (10)
$\mathrm{Label}_{\mathrm{WSI}}={\arg\max}\ \hat{p}_{c},\qquad c\in\{1,\ldots,C\},$ (11)
@@
The soft voting mechanism integrates patch classification results, facilitating the transfer of classification information from local to global levels. This weighted mechanism better accommodates the high heterogeneity issues in pathological image classification. Additionally, it simplifies the final steps of WSI classification by converting complex probability distributions into clear classification labels, providing a straightforward diagnostic basis for practical applications.
@@ 3.5.2 Training and Application of the SVM Model
Leveraging the features extracted by the patch classification model CMSwinKAN, we train a multi-class SVM model to determine whether a patch contains neuropil, Schwannian stroma, or other tissue components. During the inference phase, the SVM model receives the input feature x of each patch and outputs the corresponding category probability distribution $P_{\mathrm{SVM}}$:
$P_{\text{SVM}}=\text{softmax}\left(\begin{array}[]{c}\mathbf{w}_{0}\cdot\bm{\phi}(\mathbf{x})+b_{0},\\ \mathbf{w}_{1}\cdot\bm{\phi}(\mathbf{x})+b_{1},\\ \mathbf{w}_{2}\cdot\bm{\phi}(\mathbf{x})+b_{2}\end{array}\right),$ (12)
where the “softmax” function normalizes the SVM output values into a probability distribution. $P_{\mathrm{SVM}}=\left[P_{0},P_{1},P_{2}\right]$ represents the probabilities of the patch belonging to other tissues, neuropil, and Schwannian stroma, respectively. Based on the SVM prediction results, the voting weight of the patch $w_{\mathrm{patch}}$ is determined as follows:
@@
Table 1: Different histological types and gender composition of our private pNTs dataset.
Histological Type Total Cases (Percentage) WSI Male Female
Undifferentiated (UD) 20 (19.6%) 63 14 6
Poorly Differentiated (PD) 40 (39.2%) 56 24 16
Differentiated (D) 22 (21.6%) 62 8 14
Ganglioneuroblastoma (GNB) 10 (9.8%) 63 4 6
Ganglioneuroma (GN) 10 (9.8%) 60 3 7
Total 102 (100.0%) 304 53 49
@@ 4.1 Data Collection and Processing
The academic community faces a scarcity of large-scale, fine-grained datasets for NB due to the complexity and resource-intensive nature of data collection and annotation. This limitation restricts the availability of high-quality resources for research and development in this domain. To address this gap, we collaborated with the Department of Pediatric Oncology at Zhejiang University Children’s Hospital to establish a private dataset, Pediatric peripheral Neuroblastic Tumors(PpNTs). As shown in Table 1, this dataset includes 102 children diagnosed with NB at Zhejiang University Children’s Hospital from January 2019 to December 2023, comprising a total of 304 WSIs. Digitization was performed using a Pannoramic SCAN II digital scanner (3DHISTECH Ltd, Budapest, Hungary) at a magnification of 20×20$\times$20 ×.
@@ 4.1.1 Sliding Window Segmentation
As shown in Fig. 3, a sliding window algorithm is employed with a non-overlapping strategy. The coordinates start from the top-left corner of the WSI (0,0), and the window slides horizontally with a step size of 512 pixels. We discard these parts for the right and bottom areas that cannot be entirely covered by a $512\times 512$ pixel window, which are primarily blank regions, to avoid introducing invalid information.
4.1.2 Generation of Image Patches
Following the slicing of WSI, numerous image patches are generated, including non-tissue regions such as blank backgrounds and areas densely populated with red blood cells. These non-informative patches may introduce noise and compromise the model’s ability to effectively extract discriminative features.
@@ 4.1.3 Data Quality Improvement
To enhance data quality, this study adopts a two-stage filtering rule: White Pixel Filtering: The proportion of white pixels (RGB value $\geq$ 245) in a patch is calculated; if it exceeds 50%, the patch is deemed an invalid area such as blank background or staining defect and is discarded. Red Blood Cell Area Filtering: If the proportion of red pixels exceeds 50%, the patch is classified as a dense red blood cell area, which may obscure key pathological features and is therefore discarded.
Figure 4: Visualization of our private PpNTs dataset, which contains five types of neuroblastic tumors: Ganglioneuroma, Ganglioneuroblastoma, Poorly Differentiated Neuroblastoma, Undifferentiated Neuroblastoma, and Differentiated Neuroblastoma.
@@
4.1.4 Dataset Construction
After rigorous screening, the training and testing sets are divided based on the pathological category labels of the WSIs in an $80:20$ ratio. This results in a training set containing $2{,}379{,}858$ patches and a testing set containing $620{,}862$ patches, ensuring consistent class distribution across both datasets to avoid model performance bias due to sampling variance. Each WSI from the same case appears in only one subset to prevent data leakage. Fig. 4 provides a visual overview of our dataset, while Table 2 presents the number of WSIs and patches for each category.
@@ 4.2 Comparison on Existing Datasets
To validate the model’s generalization ability, we used the public BreakHis dataset for further testing. The BreakHis dataset is a public dataset for classifying breast cancer histopathological images collected by the P&D laboratory in Brazil. It consists of $7{,}909$ histopathological biopsy images from $82$ patients, categorized into different magnifications (×$\times 40$, ×$\times 100$, ×$\times 200$, ×$\times 400$) across a total of $8$ categories: $4$ benign types (adenosis,fibroadenoma,phyllodes tumor,tubular adenoma) and $4$ malignant types (ductal carcinoma,lobular carcinoma,mucinous carcinoma,papillary carcinoma). To thoroughly assess the generalization capability of our model(which differs the most to our own dataset in magnification), we selected images at a magnification of ×$\times 400$ and utilized the $8$ subcategories as the experimental dataset. The dataset was split into training and testing sets at an $80:20$ ratio, resulting in a certain number of WSIs. To address the class imbalance, we performed data augmentation on the training set, including brightness, contrast, and HSV color transformations, increasing the original $1{,}427$ patches to $6{,}109$ patches, while the testing set remained at $393$ patches.
@@ 4.3 Experimental Setup
The experiments were conducted on an NVIDIA Tesla V100-SXM2 GPU (32GB graphic memory) hardware platform with a software environment based on Python 3.8.20 and PyTorch 2.4.1 with CUDA 12.1 acceleration library. The training process on our private dataset lasted for 30 epochs, with a batch size of 64, taking approximately 75 hours.
For training data, input images were resized to $224\times 224$ pixels, and data augmentation techniques, such as random horizontal flipping, were applied to enhance robustness. Images were directly resized to $224\times 224$ pixels for the testing data. The model was not pre-trained; instead, model parameters were randomly initialized using a He normal distribution[36], and the AdamW optimizer was employed in conjunction with a cosine annealing algorithm. A cosine annealing strategy was utilized during training, and gradient accumulation was employed with a step count of four, effectively expanding the global batch size to 256, enhancing computational efficiency and numerical stability through AMP-mixed precision training. The model optimized for the cross-entropy loss, with a fixed batch size of 64, and implemented 20% structured pruning on the convolutional layers to reduce computational complexity.
We set $\alpha=1$ for neuropil and $\beta=8$ for Schwannian stroma, with $\gamma=1$ assigned to non-key tissue regions. This weighted fusion approach enhances diagnostic specificity by emphasizing critical tissue features while maintaining balanced contributions from all regions.
Table 4: Comparison of CMSwinKAN with other advanced models on the private PpNTs dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%) AUROC (%)
@@
VGGNet [37] 81.89 81.52 74.37 81.76 97.02
ResNet-50 [38] 92.75 92.35 90.94 92.73 99.43
DenseNet [39] 92.02 91.73 89.71 91.93 99.36
Vision Transformer [40] 89.24 89.12 85.90 89.23 98.79
Swin Transformer [41] 91.81 91.68 89.64 91.81 99.30
@@
HisImage [42] 86.29 86.26 82.78 86.21 98.00
FMDNN [26] 92.53 92.30 89.81 92.48 99.41
StoHisNet [43] 92.21 92.03 90.87 92.18 99.33
HiFuse [33] 91.49 91.40 89.31 91.49 99.29
@@
CMSwinKAN-micro 93.58 93.42 92.12 93.57 99.53
4.4 Evaluation Metrics
To comprehensively evaluate the performance of different models across various datasets, we employed the following standard evaluation metrics: Accuracy (ACC), Balanced Accuracy (BACC), Weighted F1 Score, Cohen’s Kappa Coefficient, and Area Under the Curve (AUC). The calculation of BACC requires first computing Recall and Precision.
$ACC=\frac{TP+TN}{TP+TN+FP+FN}\\ ,$ (14)
$BalancedAccuracy=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)\\ ,$ (15)
$Precision=\frac{TP}{TP+FP}\\ ,$ (16)
$Recall=\frac{TP}{TP+FN}\\ ,$ (17)
$F1=2\times\frac{{Precision}\times{Recall}}{{Precision}+{Recall}}\\ ,$ (18)
$kappa=\frac{p_{0}-p_{e}}{1-p_{e}}\\ ,$ (19)
$AUC=\frac{\sum(pi,nj)_{pi>nj}}{P\times N}.$ (20)
In the above formulas, True Positive $(TP)$ and True Negative $(TN)$ represent the samples correctly predicted as positive and negative, respectively, while False Positive $(FP)$ and False Negative $(FN)$ represent the samples incorrectly predicted as positive and negative, respectively. $P$ denotes the number of positive samples, and $N$ denotes the number of negative samples. $p_{i}$ represents the predicted score for positive samples, indicating the probability of predicting a positive sample as a positive case; $n_{j}$ represents the predicted score for negative samples, indicating the probability of predicting a negative sample as a positive case.
Table 5: Comparison of CMSwinKAN with other advanced models for WSI classification on the private PpNTs dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%)
@@ 4.4 Evaluation Metrics
CMSwinKAN-micro (soft vote) 98.41 98.33 98.01 98.39
@@ 4.5.1 Model Performance Comparison
Table 4 shows that we conducted comparative experiments on the PpNTs dataset, evaluating our model against other typical classification models, general large models in pathology, and advanced pathological image classification models. We validated the effectiveness of the models in both microscopic image classification and WSI classification. Compared to other typical models, general large models in pathology, or advanced pathological image classification models, CMSwinKAN demonstrates significant advantages in handling microscopic image classification tasks. Fig. 5 and Fig. 6 demonstrate the confusion matrices and AUC curves of CMSwinKAN tested on the PpNTs and BreakHis datasets.
Compared to the baseline models Swin Transformer and Vision Transformer, the ACC is improved by 1.77% and 4.34%, respectively. This improvement can be attributed to integrating the CMSA module and KAN in CMSwinKAN, which effectively captures fine-grained features and high-level semantic information in pathological images. This capability compensates for the limitations of the Swin Transformer in
@@
terms of field of view and lack of contextual information interaction, as well as the Vision Transformer’s shortcomings in capturing local features.
Compared with general large models in pathology, our model, which uses random initialization without pre-training, still outperformed CONCH and UNI, which were pre-trained on large-scale pathological datasets and transferred to our private dataset. While these models perform well in multimodal pathological data, their effectiveness in capturing subtle features and differentiating similar cell images in specialized fields, such as NB image classification, is limited. As illustrated in Fig. 7, CMSwinKAN demonstrates a distinct advantage in combining high accuracy with a lightweight design compared to other models.
Additionally, we compared our CMSwinKAN with advanced pathological image classification models, such as HisImage [42], FMDNN [26], and HiFuse [33], which have limitations in handling similar categorical features. Specifically, under the same training conditions, we surpassed them in all metrics; for instance, their KAPPA scores were 82.78%, 89.81%, and 89.31%, respectively, while our CMSwinKAN model achieved a KAPPA score of 92.12%. In terms of BACC, CMSwinKAN increased performance by 7.16%, 1.12%, and 2.02% over these models, respectively, significantly enhancing pathological diagnostic capabilities. Compared to the improved baseline model CTransPath based on Swin Transformer and StoHisNet using Swin Transformer Block, our method improved the ACC by 1.37%, demonstrating the architectural advantages of CMSwinKAN. In the context of WSI classification, in Table 5, the CMSwinKAN model exhibits clinical application value. After adopting a soft voting strategy, it achieved excellent performance with 98.41% ACC and 98.33% BACC, comparable to mainstream architectures like ResNet-50, DenseNet, and large pathological models like UNI. Notably, this strategy performed exceptionally well in baseline models: VGGNet’s ACC improved from 92.06% to 93.65% and BACC from 91.79% to 93.46%, confirming the beneficial impact of weighted fusion on diagnostic decisions. The model innovatively reduced the parameter count to 1.53% of that of the Swin Transformer through structural optimization, achieving an enhancement in parameter utilization efficiency while maintaining comparable WSI classification performance (98.41% ACC). This high-performance and high-efficiency characteristic allows CMSwinKAN to meet the fine-grained analysis needs of microscopic images with a KAPPA score of 92.12% and a maximum BACC improvement of 6.6% while adapting to large-scale clinical WSI processing scenarios, providing a new technical paradigm for intelligent pathological diagnosis.
@@
4.5.2 Validation of Model Generalization Ability
To further validate the generalization capability of the proposed CMSwinKAN model, we compared it with other typical classification models, large models in pathology, and advanced pathological image classification models on the publicly available breast cancer dataset, BreakHis. Table 6 shows that CMSwinKAN demonstrated strong generalization ability on the imbalanced dataset, achieving excellent results across all metrics. Fig. 8 depicts the loss and gradients on the test set during training.
Table 6: Comparison of CMSwinKAN with other advanced models on the public BreakHis dataset.
Model ACC (%) BACC (%) KAPPA (%) F1 (%) AUROC (%)
General methods
VGGNet [37] 48.09 23.57 21.44 41.37 67.71
ResNet-50 [38] 68.45 55.61 47.65 67.40 92.37
DenseNet [39] 83.21 80.81 77.33 83.64 98.25
Vision Transformer [40] 68.45 59.84 61.27 67.58 91.03
Swin Transformer [41] 75.82 71.22 68.38 76.06 96.37
Pathological large-scale pre-trained general models
UNI [22] 78.88 69.91 72.07 71.41 97.03
CTransPath [21] 83.46 86.62 78.92 83.79 98.52
CONCH [23] 86.77 86.99 86.98 86.81 94.73
Pathological classification models
HisImage [42] 64.63 55.96 52.50 57.59 84.43
FMDNN [26] 86.32 81.68 84.71 86.14 96.57
StoHisNet [43] 74.84 68.33 67.32 74.24 94.26
HiFuse [33] 62.84 47.93 48.78 60.59 87.02
Ours
CMSwinKAN-tiny 87.79 87.56 85.69 88.04 98.80
@@ 4.5.2 Validation of Model Generalization Ability
Model ACC (%) BACC (%) KAPPA (%) F1 (%)
@@
We conducted WSI classification experiments on the publicly available breast cancer histopathological dataset, BreakHis, to further confirm its generalization ability. As shown in Table 7, when using a hard voting strategy, CMSwinKAN exhibited outstanding performance even in imbalanced data scenarios: its ACC (95.06%), BACC (96.69%), KAPPA (93.40%), and F1 (95.34%) significantly surpassed all comparative models. Compared to the current optimal large pathological model, CONCH (ACC 95.05%), CMSwinKAN won by a narrow margin and led considerably by 2.94% in the BACC metric (93.75% to 96.69%); its performance improved by 1.23% compared to the classic architecture DenseNet (ACC 93.83%). This result is consistent with the conclusions drawn from the PpNTs dataset, confirming that the model effectively adapts to pathological image classification needs across different data distributions through multi-scale feature interaction with the CMSA module and nonlinear mapping in the KAN layer.

### v2

@@ 3.5.1 Voting Mechanism Design
WSI classification via hard voting assigns equal weight to each patch and determines the final result by majority rule, but its precision is relatively low. To enhance accuracy, we designed a soft voting approach that incorporates prior knowledge. In pathological analysis, patches with neuropil indicate potential malignancy (ganglioneuroblastoma, intermixed (GNBi) or NB), while patches with Schwannian stroma may indicate ganglioneuroma (GN). If Schwannian stroma comprises 50% or more of the patch, it suggests GNBi or GN; otherwise, it leans towards NB.
We can derive the voting weights for all patches in the WSI. Integrating these weights with classification results yields the final WSI classification ${\mathrm{Label}}_{\mathrm{WSI}}$:
$P\in R^{N\times C},W\in R^{N\times C},$
@@
$\hat{p}=\frac{\mathbf{s}}{\alpha},\mathrm{Label}_{\mathrm{WSI}}={\arg\max}\ \hat{p}_{c},\qquad c\in\{1,\ldots,C\},$ (10)
@@ 3.5.2 Training and Application of the SVM Model
Leveraging the features extracted by the patch classification model CMSwinKAN, we train a multi-class SVM model to determine whether a patch contains neuropil, Schwannian stroma, or other tissue components. Based on the SVM prediction results, the voting weight of the patch $w_{\mathrm{patch}}$ is determined as follows:
@@ 4.1 Data Collection and Processing
There is a shortage of pathological datasets for NB due to the complexities of data collection and annotation. To address this issue, we established a private Pediatric Peripheral Neuroblastic Tumor dataset (PpNTs), derived from 102 pediatric patients at the Children’s Hospital, Zhejiang University School of Medicine. This dataset includes three major pathological categories: NB with its histological subtypes — undifferentiated (UD), poorly differentiated (PD), and differentiated (D); GNBi; and GN. In total, 304 WSIs were acquired from these cases. Digitization was performed using a Panoramic SCAN II digital scanner (3DHISTECH Ltd, Budapest, Hungary) at a magnification of 20×. The dataset details are shown in Fig. 4, and the summary statistics are provided in Table 1. The study and the dataset was ethically approved by the Academic Ethics Committee of Children’s Hospital Zhejiang University School of Medicine (IRB No. 2023-IRB-0287-P-01). We applied for an informed consent waiver.
@@ 4.1.1 Sliding Window Segmentation
As shown in Fig. 3, a sliding window algorithm with a non-overlapping strategy is used, starting from the top-left corner of the WSI (0,0). The window slides horizontally with a step size of 512 pixels, discarding right and bottom areas that can’t be fully covered by a $512\times 512$ pixel window to avoid invalid data.
@@ 4.1.2 Data Quality Improvement
This study adopts a two-stage filtering rule: patches with over 50% white pixels (RGB $\geq$ 245) are discarded, as they likely indicate a blank background or staining defect. Patches with over 50% red pixels are classified as dense red blood cell areas and are also discarded, as they may obscure important pathological features.
Figure 4: Visual representation of our private PpNTs dataset.
@@ 4.2 Comparison on Existing Datasets
To validate the model’s generalization ability, we used the public BreakHis dataset, which contains $7{,}909$ histopathological images from $82$ patients collected by the P&D laboratory in Brazil, categorized into eight types at various magnifications. We focused on images with ×$\times 400$ magnification. To tackle class imbalance, data augmentation was applied to the training set, increasing the number of patches from $1{,}427$ to $6{,}109$, while the testing set remained at $393$ patches.
@@ 4.3 Experimental Setup
The experiments utilized an NVIDIA Tesla V100-SXM2 GPU with 32GB of graphic memory, running on Python 3.8.20 and PyTorch 2.4.1 with CUDA 12.1. Training on our private dataset lasted for 30 epochs with a batch size of 64, totaling approximately 75 hours.
Input images were resized to $224\times 224$ pixels, and the model was initialized randomly using a He normal distribution[34], and the AdamW optimizer was paired with a cosine annealing strategy.We set $\alpha=1$ for neuropil and $\beta=8$ for Schwannian stroma, with $\gamma=1$ assigned to non-key tissue regions.
Table 3: Comparison of CMSwinKAN with other advanced models on our private PpNTs and public BreakHis datasets.
Dataset PpNTs BreakHis
Model ACC BACC KAPPA F1 AUROC ACC BACC KAPPA F1 AUROC
@@
VGGNet [35] 81.89 81.52 74.37 81.76 97.02 48.09 23.57 21.44 41.37 67.71
ResNet-50 [36] 92.75 92.35 90.94 92.73 99.43 68.45 55.61 47.65 67.40 92.37
DenseNet [37] 92.02 91.73 89.71 91.93 99.36 83.21 80.81 77.33 83.64 98.25
Vision Transformer [38] 89.24 89.12 85.90 89.23 98.79 68.45 59.84 61.27 67.58 91.03
Swin Transformer [39] 91.81 91.68 89.64 91.81 99.30 75.82 71.22 68.38 76.06 96.37
@@
HisImage [40] 86.29 86.26 82.78 86.21 98.00 64.63 55.96 52.50 57.59 84.43
FMDNN [23] 92.53 92.30 89.81 92.48 99.41 86.32 81.68 84.71 86.14 96.57
StoHisNet [41] 92.21 92.03 90.87 92.18 99.33 74.84 68.33 67.32 74.24 94.26
HiFuse [30] 91.49 91.40 89.31 91.49 99.29 62.84 47.93 48.78 60.59 87.02
@@
CMSwinKAN-micro 93.58 93.42 92.12 93.57 99.53 87.79 87.56 85.69 88.04 98.80
Table 4: Comparison of CMSwinKAN with other advanced models for WSI classification on our private PpNTs dataset.
Model ACC BACC KAPPA F1
@@
CMSwinKAN-micro (soft vote) 100.00 100.00 100.00 100.00
@@ 4.4.1 Model Performance Comparison
Table 3 shows that we conducted comparative experiments on the PpNTs dataset, evaluating our model against other typical classification models, general large models in pathology, and advanced pathological image classification models. Compared to other typical models, general large models in pathology, or advanced pathological image classification models, CMSwinKAN demonstrates significant advantages in handling microscopic image classification tasks. Fig. 7 and Fig. 8 demonstrate the confusion matrices and AUC curves of CMSwinKAN tested on the PpNTs and BreakHis datasets.
@@
Compared to the baseline models Swin Transformer and Vision Transformer, the ACC is improved by 1.77% and 4.34%, respectively. This improvement can be attributed to integrating the CMSA module and KAN in CMSwinKAN, which effectively captures fine-grained features and high-level semantic information in pathological images.
Our model, using random initialization without pre-training, outperformed CONCH and UNI, which were pre-trained on large pathological datasets. As shown in Fig. 5, CMSwinKAN combines high accuracy with a lightweight design, giving it a distinct advantage over other models.
We compared our CMSwinKAN with advanced pathological image classification models like HisImage [40], FMDNN [23], and HiFuse [30], which struggle with similar categorical features. Under the same training conditions, CMSwinKAN outperformed them in all metrics, with KAPPA scores of 92.12% compared to their 82.78%, 89.81%, and 89.31%. When compared to the improved baseline model CTransPath[18] and StoHisNet[41], CMSwinKAN also improved ACC by 1.37%, highlighting its architectural advantages. For WSI classification, the CMSwinKAN model demonstrates significant clinical value. With a soft voting strategy, it achieved 98.41% ACC and 98.33% BACC. This strategy notably improved VGGNet’s performance, raising ACC from 92.06% to 93.65% and BACC from 91.79% to 93.46%.
4.4.2 Validation of Model Generalization Ability
To further validate the generalization capability of the proposed CMSwinKAN model, we compared it with other typical classification models, large models in pathology, and advanced pathological image classification models on the publicly available breast cancer dataset, BreakHis. Table 3 with BreakHis dataset shows that CMSwinKAN demonstrated strong generalization ability on the imbalanced dataset, achieving excellent results across all metrics. Fig. 6 depicts the loss and gradients on the test set during training.
@@ 4.4.2 Validation of Model Generalization Ability
Model ACC BACC KAPPA F1
@@
We conducted WSI classification experiments on the BreakHis dataset to further confirm its generalization ability. As shown in Table 5, when using a hard voting strategy, CMSwinKAN exhibited outstanding performance even in imbalanced data scenarios. This result aligns with findings from the PpNTs dataset, confirming that the model adapts well to pathological image classification across various data distributions through multi-scale feature interaction in the CMSA module and nonlinear mapping in the KAN layer.